OpenStudy (anonymous):

which equation represent y=x^2-10x+30 in vertex form

OpenStudy (amistre64):

which which equation ?

OpenStudy (amistre64):

the one that looks like: y = M(x-xo)^2 +yo

OpenStudy (amistre64):

(x-5)^2 +30-25 if i see it right

OpenStudy (amistre64):

y = (x-5)^2 +5 .... just complete the square; and remember to subtract the part you add to keep theing copacetic

OpenStudy (anonymous):

a. y=(x-5)^2+5 b. y=(x-5)^2-5 c. y=(x-5)^2+30 d. y=(x-5)^2+55

OpenStudy (amistre64):

i believe its 'a'; you can dbl check by expanding it out again